steve is 5 times as as old as janice in 12 years she will be twice as old as she was then, how old are they now?
Answer by mananth(16949)

steve is 5 times as as old as janice in 12 years he will be twice as old as she was then, how old are they now
I have corrected the sum . I do not know if the correction is OK.
let janice be x years now
steve will be5x
after 12 years
5x +12 = 2(2x+12)
5x+12 = 4x +24
5x-4x= 24-12
x=12 janice's age now
Steve's age will be 12* 5=60 years